Use the table to answer the questions. Assume firms pay all profits out to resource owners, there is no depreciation, and there are no taxes. 2010 (C) $350 $100 $175 $75 $50 Government spending (G) $125 $10 Rent Profit Investment spending ) Interest Net exports (NX) Employee compensation Enter the value of GDP in 2010. Use the factor income approach to calculate employee compensation in 2010.
Answer #1

1.

GDP in 2010=C+I+G+NX

=350+75+125+10

=$560

2.

Income approach

GDP = Compensation of employees+ rent+ interest+ profit+ Mixed income of self-employed+ depreciation

560= Compensation of employees +100+50+175+0+0

560= Compensation of employees+325

Compensation of employees=560-325

=$235